OpenStudy (anonymous):

please help......find f '(a)...... f(t)=(2t+1)/(t+3)

OpenStudy (cwrw238):

using the quotient rule: f'(x) = 2(t+3) - (2t+1) ------------- (t+3)^2 = 5 ----- (t+3)^2 f'(a) = 5 / (a + 3)^2

OpenStudy (anonymous):

thank you so much , i got it

OpenStudy (cwrw238):

yw